# Bi-Lipschitz Smoothing under Ricci and Injectivity Bounds

### v1

@@ 2 Main result
so $|v|_{h}\leq C^{1/2}|v|_{g}$ and $|v|_{h}\geq C^{-1/2}|v|_{g}$. We now integrate along the curves and obtain the required distance bounds. For the volume forms, we choose a $g$-orthonormal basis so that $g$ is represented by $I$ and $h$ by a positive definite matrix $A$. It follows that $C^{-1}I\leq A\leq CI$, so we infer that $\det A\in[C^{-n},C^{n}]$ and $d\operatorname{vol}_{h}=\sqrt{\det A}\,d\operatorname{vol}_{g}$. We obtain completeness directly from the distance comparison. Observe that an $h$-Cauchy sequence is $g$-Cauchy, so it converges in $(M,g)$. In particular, it also converges in $(M,h)$ by the reverse comparison, which concludes the proof. ∎
@@
We will now consider a conjugate-radius lower bound. To this end, for $p\in M$, let $\operatorname{conj\,rad}_{\bar{g}}(p)$ denote the conjugate radius at $p$. Observe that if $\exp_{p}$ is a diffeomorphism on the open $\bar{g}$-ball of radius $r$ in $T_{p}M$, then $\exp_{p}$ has no critical points there. This implies that $r\leq\operatorname{conj\,rad}_{\bar{g}}(p)$. We then obtain:
@@
and such that its weak $C^{2,\alpha}$-norm on scale $1$ is bounded:
@@
where $Q_{2}$ depends only on $n$, $\alpha$, $\varepsilon$ (here $1$), $m=2$, and $Q_{0}(1)$.

### v2

@@ 2 Main result
so $|v|_{h}\leq C^{1/2}|v|_{g}$ and $|v|_{h}\geq C^{-1/2}|v|_{g}$. We now integrate along the curves and obtain the required distance bounds. For the volume forms, we choose a $g$-orthonormal basis so that $g$ is represented by II and $h$ by a positive definite matrix AA. It follows that $C^{-1}I\leq A\leq CI$, so we infer that $\det A\in[C^{-n},C^{n}]$ and $d\operatorname{vol}_{h}=\sqrt{\det A}\,d\operatorname{vol}_{g}$. We obtain completeness directly from the distance comparison. Observe that an $h$-Cauchy sequence is $g$-Cauchy, so it converges in $(M,g)$. In particular, it also converges in $(M,h)$ by the reverse comparison, which concludes the proof. ∎
@@
We will now consider a conjugate-radius lower bound. To this end, for $p\in M$, let $\operatorname{conj\,rad}_{\bar{g}}(p)$ denote the conjugate radius at pp. Observe that if $\exp_{p}$ is a diffeomorphism on the open $\bar{g}$-ball of radius $r$ in $T_{p}M$, then $\exp_{p}$ has no critical points there. This implies that $r\leq\operatorname{conj\,rad}_{\bar{g}}(p)$. We then obtain:
@@
and such that its weak $C^{2,\alpha}$-norm on scale 11 is bounded:
@@
where $Q_{2}$ depends only on $n$, $\alpha$, $\varepsilon$ (here 11), $m=2$, and $Q_{0}(1)$.